Know how to calculate how long it will take an amoeba to reach a certain mass...

  1. Know how to calculate how long it will take an amoeba to reach a certain mass (in this questions, it was the mass of an elephant, which we had to guess ourselves), we were given that the doubling time for an amoeba is 23 hours. Also, remember that the size of an amoeba is the size of an average eukaryotic cell, not a bacterial cell.
  2. Find number of proteins in an amoeba, knowing that the concentration of proteins in an amoeba is 200mg/ml. Remember, again, that an amoeba is a eukaryote.
  3. Describe, starting from mRNA outside the nucleus, how a 7 folded transmembrane protein ends up on the plasma membrane.
Answer #1

The average mass of an elephant= 5*10^6 grams

1 amoeba weighs = 0.00018 grams

Therefore, doubling time = 23 hours

A=A0 * (1/2)^ t/h

5*10^6= 0.00018 * (1/2)^t/23

t= 23/10 * 10^10 hours

It depends on the type of amoeba and its weight. Cannot calculate directly

The proteins, however, are made as single-pass transmembrane proteins in the ER. While still in the ER, the transmembrane segment of the protein is cleaved off and a glycosylphosphatidylinositol (GPI) anchor is added.